Carbohydrates are an important class of organic molecules that act as energy storage compounds in living systems. A 0.8361g sample of a particular carbohydrate is analysed and found to contain 0.4565g of carbon and 0.0761g of hydrogen, with the rest of the mass being made up of oxygen. The molar mass of the compound is 88g mol−1. Determine both the empirical and molecular formulae of the compound
1
Expert's answer
2020-05-05T04:01:44-0400

0.8361-0.4565-0.0761=0.3035 g of oxygen.

0.4565/12: 0.0761/1: 0.3035/16 =(0.038:0.0761:0.0189)/0.0189= 2:4:1.

2 atom of C, 4 atom of H, 1 atom of O.

Summ = 44.

88/44= 2.

C4H8O2.